A particle undergoes SHM with a time period of 2 seconds in how much time will it travel from its mean position to a displacement equal to half of its amplitude
a-0.5 sec
b-1/3sec
c-1/4sec
d-1/6sec

Grade:12

1 Answers

Sunil Kumar FP
10 years ago
x=Asine omega t
x=A sin 2pi/T *t
A/2=A sin 2pi/T *t
2pi/T *t=pi/6
t=1/6 s
thanks and regards
sunil kr
askIITian faculty